# Automated Federated Pipeline for Parameter-Efficient Fine-Tuning of Large Language Models

### v1

@@ 2.3. Weight-Level Configuration
Although LoRA has reduced large computing overhead (see Section 2.1) for LLMs, due to the constrained computing capacity at an edge server, training all weights with LoRA in each layer still poses a heavy workload. We train multiple LoRA adapter versions by increasing the number of trainable weights in each layer shown in Figure LABEL:fig:diff_config:layers. The perplexity (PPL) is used as the performance indicator to evaluate how well the model has learned the distribution of the data it was trained on, and a smaller PPL represents better performance. Apparently, the more weights the LoRA adapter has, the higher the computing overhead, and the better the performance required and obtained. A naive solution is to configure the number of trainable weights based on a computing budget.
@@
However, how to find the prior weights is a non-trivial task. We have conducted another measurement experiment to study the impact of weight indices in LoRA training. We select four trainable weight vectors (i.e., $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$, $\mathbf{W}_{o}$ shown in Figure 2) from the transformer block to train LoRA adapters in an round illustrated in Figure LABEL:fig:diff_config:chan. It demonstrates that various weight indices have different performances. We also plot PLLs with four trainable weight vectors in different training rounds shown in Figure LABEL:fig:important_ppl_20round, and calculate their standard deviations (STDs) illustrated in Figure LABEL:fig:important_std. Apparently, the importance of four trainable weight vectors changes for the LoRA adapter in each training round, and hence, a one-shot measurement cannot identify their priority. Therefore, we need to investigate how to identify and prioritize the important trainable weights automatically for each edge server under its computing budget.
2.4. Budget-Aware Model Parameters Alignment
@@
Figure 7. Performance and computing consumption of GPT-2 with LoRA adapter with different training batch sizes and ranks.
@@
While the training weights are determined, several parameters still impact LoRA training, in particular, batch size, decomposition rank, dropout, alpha, etc. We evaluate all of them and find that batch size and decomposition rank are the two most important parameters for training performance [^1] The same results are also found in (Sheng et al., 2023; Li et al., 2024). To study the impact of both batch size and decomposition rank, we increase them, respectively, to train the pre-trained GPT-2 with LoRA, and the results are illustrated in Figure 7. It is clear to observe that a larger batch size or decomposition rank always offers better performance but consumes more computing resources. However, given the computing budget of an edge server, we cannot always select the largest batch size and rank. We need to find an optimal combination of batch size and rank under the computing budget for each edge server in an automated federated pipeline.
3. Automated Federated Pipeline Design
3.1. Overview of FedPipe
@@
To tackle the above challenges, we design FedPipe, an automated federated pipeline shown in Figure 8, to determine the optimal fine-tuning model structure to improve the performance of fine-tuning LLMs, accommodating heterogeneous and resource-constrained edge servers. We first theoretically formulate an MILP to model the straggler problem under heterogeneous and constrained computing resources in LLM FL to guide FedPipe’s design (Section 3.2). To solve the problem, we propose two-level solutions to identify the important weights (Section 3.3) and heterogeneous adapters’ configuration (Section 3.4). We also consider heterogeneous storage memories and quantize the LLM based on the budget of an edge server (Section 3.5). Finally, we design an effective method to aggregate all local models with different size of LoRA adapters (Section 3.6). After that, the next training round starts.
Figure 8. The system overview of FedPipe.
3.2. Modelling Heterogeneous LoRA Adapters
@@
In this section, we model the LoRA adapters for different edge servers with heterogeneous computing resources. As illustrated in Figure 9, $N$ edge servers independently train LoRA adapters, each with its computing budget $C_{i,\max}$ for the $i$-th edge server. After training, the central server aggregates all LoRA adapters of edge servers. LoRA is built on the insights that over-parametrized models essentially reside on a low intrinsic dimension (Aghajanyan et al., 2021), and thus, can be characterized with the reduced number of trainable parameters. LoRA augments the parameters of the frozen pre-trained model with an additional factorized projection, which constrains weight updates by representing them with a low-rank decomposition $\mathbf{W}=\mathbf{W}_{0}+\Delta\mathbf{W}$, where $\Delta\mathbf{W}=\mathbf{BA}$, $\mathbf{B}\in\mathbf{R}^{d_{i}\times r}$ and $\mathbf{A}\in\mathbf{R}^{r\times d_{o}}$ denote the incremental matrix and rank decomposition matrices, respectively. The rank is $r\ll\min{(d_{i},d_{o})}$ (e.g., $r=8$ when $d_{i}=d_{o}=1024$), and the dimensions of input and output of rank decomposition matrices $\bf B$ and $\bf A$ are $d_{i}$ and $d_{o}$, respectively. The LoRA adapter is the cascade structure composed of $\mathbf{B}$ and $\mathbf{A}$.
Figure 9. Integrating LoRA adapters into LLM FL.
@@
The trainable weight of the $i$-th edge server is denoted by $\mathbf{W}_{i}$, and its local training dataset $\mathcal{D}_{i}=\left\{\mathbf{x}_{i,j},y_{i,j}\right\}$, where $\mathbf{x}_{i,j}$ and $y_{i,j}$ represent the $j$-th input data and its corresponding label, respectively. Thus, the local loss function for the $i$-th edge server can be denoted as $L_{i}(\mathbf{W}_{i})=\frac{1}{\left|\mathcal{D}_{i}\right|}\sum_{j=1}^{\left|\mathcal{D}_{i}\right|}L_{i,j}(\mathbf{x}_{i,j},y_{i,j};\mathbf{W}_{i})$ where $L_{i,j}(\mathbf{x}_{i,j},y_{i,j};\mathbf{W}_{i})$ represents the sample-wise loss function for the $j$-th data sample in the local dataset $\mathcal{D}_{i}$. Denote the total dataset as $\mathcal{D}={\textstyle\bigcup_{i=1}^{N}\mathcal{D}_{i}}$, the global loss function is expressed as the weighted average of the local loss functions. To find the optimal model parameters $\mathbf{W}^{\ast}$, the learning objective of the global model could be formulated as
(1) $\min_{\mathbf{W}}{L(\mathbf{W})}=\min_{\mathbf{W}}\sum_{i=1}^{N}\frac{\left|\mathcal{D}_{i}\right|}{\left|\mathcal{D}\right|}{L_{i}(\mathbf{W})}$
@@
Recalling Section 2.3, batch size and rank affect the LoRA training performance directly, and thus, we rewrite Eqn. (1) in the following way
(2a) $\min_{b_{i},\mathcal{R}_{i},\mathcal{W}_{i}}\quad$ $\sum_{i=1}^{N}\frac{\left|\mathcal{D}_{i}\right|}{\left|\mathcal{D}\right|}{L_{i}(\mathbf{W}|b_{i},\mathcal{R}_{i},\mathcal{W}_{i})},$
(2b) s.t. $b_{min}\leq b_{i}\leq b_{max},$
(2c) $\mathcal{W}_{i}\in\{\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}\},$
(2d) ${r}_{i}^{m}\in\mathcal{Q}^{m},$
(2e) ${b_{i}\cdot f_{c}(\mathcal{W}_{i},\mathcal{R}_{i})\leq C_{i,\max}},$
@@
where $\mathcal{W}_{i}$ and $\mathcal{R}_{i}=\left\{{r}_{i}^{1},{r}_{i}^{2},...,{r}_{i}^{M}\right\}$ represent the sets of selected trainable weights and corresponding ranks, constraint Eqn. (2b) refers to the range of the batch size for model training, and $b_{i}$ is the batch size of the $i$-th edge server. The constraint Eqn. (2c) is used to pick up important weights in the transformer layer (see Section 2.3). Eqn. (2d) guarantees that the rank of the $m$-th weight is selected in $\mathcal{Q}^{m}$ where $m$ ranges from 1 to $M=4$ to represent different weights (i.e., $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$). Eqn. (2e) indicates the computing budget for the $i$-th edge server where $f_{c}(\cdot)$ maps the relationship between trainable weight structure to floating point operations per second (FLOPS) (Kaplan et al., 2020). The optimization problem (3.2) is an MILP problem, which is typically NP-hard. Hence, we propose an automated federated pipeline framework to solve this problem from Section 3.3 to 3.5.
Figure 10. The average normalized important metric and PPL of different trainable weights for GPT-2 models across 20 training rounds.
3.3. Identification of important weights
@@
Recalling in Figure 5 (see Section 2.3), when the whole training process is done, different weights have various average performances on the LoRA adapter. In practice, the situation is more complex, since different combinations of weights also result in various training performances. More importantly, under a computing budget, the $i$-th edge server needs to identify and predict the combinations of weights with LoRA adapters for the next round of training.
@@
To identify a set of important weights for each edge server, FedPipe introduces an indicator for important weights to quantify the important level of parameters and the sensitivity of parameters in the training loss. The well-known principal components analysis (PCA) (Labrín and Urdinez, 2020) algorithm identifies the most information-rich dimensions with the larger singular value by employing matrix singular value decomposition (SVD). Motivated by this, we perform the SVD on $\mathbf{BA}$ (see Section  3.2) of each trainable weight before each round of the training.
@@
Since larger singular values indicate richer data information, we characterize the importance of weights using the average singular value of each trainable weight. Additionally, we also devise a sensitivity-based importance scoring function based on sensitivity-based importance scoring (Molchanov et al., 2019; Zhang et al., 2022; Liang et al., 2021). Therefore, at the $t$-th training round, to properly measure the parameter contribution of the $m$-th trainable weight (LoRA adapter) for $i$-th edge server, the importance metric can be calculated as follows
(3) ${s^{m{\kern 1.0pt},t}_{i}}=\frac{1}{{{d_{1}}}}\sum\limits_{j=1}^{{d_{1}}}{{\lambda^{m{\kern 1.0pt},t}_{i,j}}}+{\phi}\left({{{\bf{W}}^{m{\kern 1.0pt},t}_{i}}}\right)$
@@
(4) $I\left({\bf{W}}\right)=\frac{1}{d_{2}}\sum\limits_{j=1}^{{d_{2}}}{\left|{{w_{j}}{\nabla_{{w_{j}}}}L\left({\bf{W}}\right)}\right|}$
where ${w_{j}}$ and ${{\nabla_{{w_{j}}}}L\left({\bf{W}}\right)}$ represent the $j$-th trainable parameter of the model parameter ${\bf{W}}$ and its corresponding gradient, respectively, and $d_{2}$ is the total number of model parameter $\bf W$.
Eqn. (4) essentially approximates the average change in loss per parameter when a parameter is zeroed out (Molchanov et al., 2019; Liang et al., 2021). However, the sensitivity in Eqn. (4) is not a reliable indicator of the importance of ${\bf W}$. This is because it is estimated based on one mini-batch data sample, where random sampling and complex training dynamics result in high variability and significant uncertainty in sensitivity estimation using Eqn. (4). Therefore, at the $t$-th training round, we address this problem by designing sensitivity smoothing and uncertainty quantification:
(5) ${{\overline{I}}^{t}}\left({\bf{W}}\right)={\lambda_{1}}{{\overline{I}}^{t-1}}\left({\bf{W}}\right)+\left({1-{\lambda_{1}}}\right){I}^{t}\left({\bf{W}}\right)$
(6) ${{\overline{U}}^{t}}\left({\bf{W}}\right)={\lambda_{2}}{{\overline{U}}^{t-1}}\left({\bf{W}}\right)+\left({1-{\lambda_{2}}}\right)\left|{{I}^{t}\left({\bf{W}}\right)-{{\overline{I}}^{t}}\left({\bf{W}}\right)}\right|$
where $10<{\lambda_{1}},{\lambda_{2}}<$ 1, ${\overline{I}}\left({{{\cdot}}}\right)$ represents the smoothed sensitivity by exponential moving average, and ${\overline{U}}\left({{{\cdot}}}\right)$ denotes the uncertainty term. Subsequently, the sensitivity-based importance scoring function ${\phi}\left({{{\cdot}}}\right)$ is defined as the product of ${\overline{I}}\left({{{\cdot}}}\right)$ and ${\overline{U}}\left({{{\cdot}}}\right)$:
(7) ${\phi}^{t}\left({\bf{W}}\right)={{\overline{I}}^{t}}\left({\bf{W}}\right){{\overline{U}}^{t}}\left({\bf{W}}\right)$
Figure 11. The automated LoRA adapters configuration solution.
As shown in Figure 10, we observe that trainable weights with higher importance metrics exhibit better fine-tuning performance, i.e., lower PPL values, which further illustrates the effectiveness of the proposed importance metrics. With the importance of weights $\mathbf{s}_{i}=\{s^{m}_{i}$}, the $i$-th edge server can configure its set of LoRA adapters in the following section.
3.4. Heterogeneous LoRA Adapters Configuration
We propose an automatic configuration solution to select batch size and rank for each edge server, based on the proposed important weight indicators. Since exhaustive search for all kinds of batch sizes, weights, and ranks is impossible due to the large search space, FedPipe leverages a two-stage search algorithm to tackle this problem as shown in Figure 11. The first stage is to determine batch size, according to data diversity, and the second stage is to find the suitable rank under a computing budget.

### v2

@@ II-C Weight-Level Configuration
Although LoRA has reduced large computing overhead (see Section II-A) for LLMs, due to the constrained computing capacity at an edge server, training all weights with LoRA in each layer still poses a heavy workload. We train multiple LoRA adapter versions by increasing the number of trainable weights in each layer shown in Fig. 5a. The perplexity (PPL) is used as the performance indicator to evaluate how well the model has learned the distribution of the data it was trained on, and a smaller PPL represents better performance. Apparently, the more weights the LoRA adapter has, the higher the computing overhead, and the better the performance required and obtained. A naive solution is to configure the number of trainable weights based on a computing budget.
@@
However, how to find the prior weights is a non-trivial task. We have conducted another measurement experiment to study the impact of weight indices in LoRA training. We select four trainable weight vectors (i.e., $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$, $\mathbf{W}_{o}$ shown in Fig. 2) from the transformer block to train LoRA adapters in an round illustrated in Fig. 5b. It demonstrates that various weight indices have different performances. We also plot PLLs with four trainable weight vectors in different training rounds shown in Fig. 6a, and calculate their standard deviations (STDs) illustrated in Fig. 6b. Apparently, the importance of four trainable weight vectors changes for the LoRA adapter in each training round, and hence, a one-shot measurement cannot identify their priority. Therefore, we need to investigate how to identify and prioritize the important trainable weights automatically for each edge server under its computing budget.
II-D Budget-Aware Model Parameters Alignment
(a) Batch size
(b) Decomposition rank
@@ II-D Budget-Aware Model Parameters Alignment
Figure 7: Performance and computing consumption of GPT-2 with LoRA adapter with different training batch sizes and ranks.
@@
While the training weights are determined, several parameters still impact LoRA training, in particular, batch size, decomposition rank, dropout, alpha, etc. We evaluate all of them and find that batch size and decomposition rank are the two most important parameters for training performance [^1] The same results are also observed in [42, 54]. To study the impact of both batch size and decomposition rank, we increase them, respectively, to train the pre-trained GPT-2 with LoRA, and the results are illustrated in Fig. 7. It is clear to observe that a larger batch size or decomposition rank always offers better performance but consumes more computing resources. However, given the computing budget of an edge server, we cannot always select the largest batch size and rank. We need to find an optimal combination of batch size and rank under the computing budget for each edge server in an automated federated pipeline.
III Automated Federated Pipeline Design
III-A Overview of FedPipe
@@
To tackle the above challenges, we design FedPipe, an automated federated pipeline shown in Fig. 8, to determine the optimal fine-tuning model structure to improve the performance of fine-tuning LLMs, accommodating heterogeneous and resource-constrained edge servers. We first theoretically formulate an MILP to address the straggler problem under heterogeneous and constrained computing resources in LLM FL to guide FedPipe’s design (Section III-B). To solve the problem, we propose two-level solutions to identify the important weights (Section III-C) and heterogeneous adapters’ configuration (Section III-D). We also consider heterogeneous storage memories and quantize the LLM based on the budget of an edge server (Section III-E). Finally, we design an effective method to aggregate all local models with different size of LoRA adapters (Section III-F). After that, the next training round starts.
Figure 8: The system overview of FedPipe.
III-B Modelling Heterogeneous LoRA Adapters
@@
In this section, we model the LoRA adapters for different edge servers with heterogeneous computing resources. As illustrated in Fig. 9, $N$ edge servers independently train LoRA adapters, each with its computing budget $C_{i,\max}$ for the $i$-th edge server. After training, the central server aggregates all LoRA adapters of edge servers. LoRA is built on the insights that over-parametrized models essentially reside on a low intrinsic dimension [55], and thus, can be characterized with the reduced number of trainable parameters. LoRA augments the parameters of the frozen pre-trained model with an additional factorized projection, which constrains weight updates by representing them with a low-rank decomposition $\mathbf{W}=\mathbf{W}_{0}+\Delta\mathbf{W}$, where $\Delta\mathbf{W}=\mathbf{BA}$, $\mathbf{B}\in\mathbf{R}^{d_{i}\times r}$ and $\mathbf{A}\in\mathbf{R}^{r\times d_{o}}$ denote the incremental matrix and rank decomposition matrices, respectively. The rank is $r\ll\min{(d_{i},d_{o})}$ (e.g., $r=8$ when $d_{i}=d_{o}=1024$), and the dimensions of input and output of rank decomposition matrices $\bf B$ and $\bf A$ are $d_{i}$ and $d_{o}$, respectively. The LoRA adapter is the cascade structure composed of $\mathbf{B}$ and $\mathbf{A}$.
@@
The trainable weight of the $i$-th edge server is denoted by $\mathbf{W}_{i}$, and its local training dataset $\mathcal{D}_{i}=\left\{\mathbf{x}_{i,j},y_{i,j}\right\}$, where $\mathbf{x}_{i,j}$ and $y_{i,j}$ represent the $j$-th input data and its corresponding label, respectively. Thus, the local loss function for the $i$-th edge server can be denoted as $L_{i}(\mathbf{W}_{i})=\frac{1}{\left|\mathcal{D}_{i}\right|}\sum_{j=1}^{\left|\mathcal{D}_{i}\right|}L_{i,j}(\mathbf{x}_{i,j},y_{i,j};\mathbf{W}_{i})$ where $L_{i,j}(\mathbf{x}_{i,j},y_{i,j};\mathbf{W}_{i})$ represents the sample-wise loss function for the $j$-th data sample in the local dataset $\mathcal{D}_{i}$. Denote the total dataset as $\mathcal{D}={\textstyle\bigcup_{i=1}^{N}\mathcal{D}_{i}}$, the global loss function is expressed as the weighted average of the local loss functions. To find the optimal model parameters $\mathbf{W}^{\ast}$, the learning objective of the global model can be formulated as
$\min_{\mathbf{W}}{L(\mathbf{W})}=\min_{\mathbf{W}}\sum_{i=1}^{N}\frac{\left|\mathcal{D}_{i}\right|}{\left|\mathcal{D}\right|}{L_{i}(\mathbf{W})}$ (1)
@@
Recalling Section II-C, batch size and rank affect the LoRA training performance directly, and thus, we rewrite Eqn. (1) in the following way
$\min_{b_{i},\mathcal{R}_{i},\mathcal{W}_{i}}\quad$ $\sum_{i=1}^{N}\frac{\left|\mathcal{D}_{i}\right|}{\left|\mathcal{D}\right|}{L_{i}(\mathbf{W}|b_{i},\mathcal{R}_{i},\mathcal{W}_{i})},$ (2a)
s.t. $b_{min}\leq b_{i}\leq b_{max},$ (2b)
$\mathcal{W}_{i}\in\{\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}\},$ (2c)
${r}_{i}^{m}\in\mathcal{Q}^{m},$ (2d)
${b_{i}\cdot f_{c}(\mathcal{W}_{i},\mathcal{R}_{i})\leq C_{i,\max}},$ (2e)
@@
where $\mathcal{W}_{i}$ and $\mathcal{R}_{i}=\left\{{r}_{i}^{1},{r}_{i}^{2},...,{r}_{i}^{M}\right\}$ represent the sets of selected trainable weights and corresponding ranks, constraint Eqn. (2b) refers to the range of the batch size for model training, and $b_{i}$ is the batch size of the $i$-th edge server. The constraint Eqn. (2c) is used to pick up important weights in the transformer layer (see Section II-C). Eqn. (2d) guarantees that the rank of the $m$-th weight is selected in $\mathcal{Q}^{m}$ where $m$ ranges from 1 to $M=4$ to represent different weights (i.e., $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$). Eqn. (2e) indicates the computing budget for the $i$-th edge server where $f_{c}(\cdot)$ maps the relationship between trainable weight structure to floating point operations per second (FLOPS) [56]. The optimization problem (2) is an MILP problem, which is typically NP-hard. Hence, we propose an automated federated pipeline framework to solve this problem from Section III-C to III-E.
Figure 9: Integrating LoRA adapters into LLM FL.
III-C Identification of important weights
@@
Recalling in Fig. 5 (see Section II-C), when the whole training process is done, different weights have various average performances on the LoRA adapter. In practice, the situation is more complex, since different combinations of weights also result in various training performances. More importantly, under a computing budget, the $i$-th edge server needs to identify and predict the combinations of weights with LoRA adapters for the next round of training.
@@
To identify a set of important weights for each edge server, FedPipe introduces an indicator for important weights to quantify the important level of parameters and the sensitivity of parameters in the training loss. The well-known principal components analysis (PCA) [57] algorithm identifies the most information-rich dimensions with larger singular values by employing matrix singular value decomposition (SVD). Motivated by this, we perform the SVD on $\mathbf{BA}$ (see Section  III-B) of each trainable weight before each round of the training.
@@
Since larger singular values contain richer data information, we characterize the importance of weights using the average singular value of each trainable weight. Additionally, we also devise a sensitivity-based importance scoring function based on sensitivity-based importance scoring [58, 59, 60]. Therefore, at the $t$-th training round, to properly measure the parameter contribution of the $m$-th trainable weight (LoRA adapter) for $i$-th edge server, the importance metric can be calculated as follows
${s^{m{\kern 1.0pt},t}_{i}}=\frac{1}{{{d_{1}}}}\sum\limits_{j=1}^{{d_{1}}}{{\lambda^{m{\kern 1.0pt},t}_{i,j}}}+{\phi}\left({{{\bf{W}}^{m{\kern 1.0pt},t}_{i}}}\right)$ (3)
@@
$I\left({\bf{W}}\right)=\frac{1}{d_{2}}\sum\limits_{j=1}^{{d_{2}}}{\left|{{w_{j}}{\nabla_{{w_{j}}}}L\left({\bf{W}}\right)}\right|}$ (4)
where ${w_{j}}$ and ${{\nabla_{{w_{j}}}}L\left({\bf{W}}\right)}$ represent the $j$-th trainable parameter of the model parameter ${\bf{W}}$ and its corresponding gradient, respectively. $d_{2}$ is the total number of model parameter $\bf W$.
Eqn. (4) essentially approximates the average change in loss per parameter when a parameter is zeroed out [58, 60]. However, the sensitivity in Eqn. (4) is not a reliable indicator of the importance of ${\bf W}$. This is because it is estimated based on one mini-batch data sample, where random sampling and complex training dynamics result in high variability and significant uncertainty in sensitivity estimation using Eqn. (4). Therefore, at the $t$-th training round, we address this problem by designing sensitivity smoothing and uncertainty quantification [59]:
${{\overline{I}}^{t}}\left({\bf{W}}\right)={\lambda_{1}}{{\overline{I}}^{t-1}}\left({\bf{W}}\right)+\left({1-{\lambda_{1}}}\right){I}^{t}\left({\bf{W}}\right)$ (5)
${{\overline{U}}^{t}}\left({\bf{W}}\right)={\lambda_{2}}{{\overline{U}}^{t-1}}\left({\bf{W}}\right)+\left({1-{\lambda_{2}}}\right)\left|{{I}^{t}\left({\bf{W}}\right)-{{\overline{I}}^{t}}\left({\bf{W}}\right)}\right|$ (6)
where $10<{\lambda_{1}},{\lambda_{2}}<$ 1, ${\overline{I}}\left({{{\cdot}}}\right)$ represents the smoothed sensitivity by exponential moving average, and ${\overline{U}}\left({{{\cdot}}}\right)$ denotes the uncertainty term. Subsequently, the sensitivity-based importance scoring function ${\phi}\left({{{\cdot}}}\right)$ is defined as the product of ${\overline{I}}\left({{{\cdot}}}\right)$ and ${\overline{U}}\left({{{\cdot}}}\right)$ [59]:
${\phi}^{t}\left({\bf{W}}\right)={{\overline{I}}^{t}}\left({\bf{W}}\right){{\overline{U}}^{t}}\left({\bf{W}}\right)$ (7)
(a) Normalized important metric.
(b) PPL
Figure 10: The average normalized important metric and PPL of different trainable weights for GPT-2 models across 20 training rounds.
As shown in Fig. 10, we observe that trainable weights with higher importance metrics exhibit better fine-tuning performance, i.e., lower PPL values, which further illustrates the effectiveness of the proposed importance metrics. With the importance of weights $\mathbf{s}_{i}=\{s^{m}_{i}$}, the $i$-th edge server can configure its set of LoRA adapters in the following section.
Figure 11: The automated LoRA adapters configuration solution.
III-D Heterogeneous LoRA Adapters Configuration
We propose an automatic configuration solution to select batch size and rank for each edge server, based on the proposed important weight indicators. Since exhaustive search for all kinds of batch sizes, weights, and ranks is impossible due to the large search space, FedPipe leverages a two-stage search algorithm to tackle this problem as shown in Fig. 11. The first stage is to determine batch size, according to data diversity, and the second stage is to find the suitable rank under a computing budget.